# FedBaF: Federated Learning Aggregation Biased by a Foundation Model

### v1

@@ 1 Introduction
Developing foundation models [1] has become a major focus for leading technology companies like OpenAI, Microsoft, and Amazon AWS. These deep learning models are often trained with vast amounts of high-quality data [2] and their ability to generalize across different tasks and domains has made them essential assets for industry, government, and academia. Foundation models have been applied to natural language processing (e.g., text generation, translation, summarization), image generation and recognition, healthcare diagnostics, finance predictive analytics, and customer service and virtual assistant tasks [3, 4, 5, 6, 7]. When a foundation model’s training data distribution overlaps with a new application, it provides a robust starting point for fine-tuning and customization. Instead of training a model from scratch, with limited data and compute, we can leverage pre-trained foundation models to enable faster training.
For many applications, data that could be used to customize or fine-tune foundation models is often distributed across multiple clients, such as a network of clinics or small companies spread across different jurisdictions. For example, fine-tuning a recommendation model to fit a small company’s product offering may require data from clients in various regions; and adapting a healthcare model for a network of clinics would involve distributed, confidential data sources. Therefore, effective generalization requires access to diverse data from multiple clients with similar objectives [8].
@@
Federated Learning (FL) is a promising solution for fine-tuning these models without sharing client data: FL clients train models on diverse local data and a central FL server aggregates the client updates to build and refine a global model [9, 10, 11, 12, 13, 14]. Using a foundation model to initialize the FL global model leads to effective customization that leverages diverse, distributed data without directly accessing the client data [15, 3]. However, there are significant risks associated with sending a foundation model to clients, as required in traditional FL fine-tuning methods.
First, disclosing a foundation model’s weights to FL clients poses a significant security risk. For example, malicious actors could carry out membership inference attacks, identifying whether specific data was part of the foundation model’s training dataset. Then, an attacker could disrupt the global model’s training by introducing updates that degrade performance on the identified data [16, 17, 18]. Similarly, for model inversion attacks, attackers reverse-engineer the model weights to infer sensitive training data (see Figure 1[19, 20, 21]. Protecting foundation models, often trained on sensitive, proprietary data, is critical for safeguarding the training data and maintaining model integrity [22, 23].
Second, in competitive business contexts, disclosing foundation model weights to clients risks leaking strategic insights and proprietary information to adversaries [4, 24]. This undermines a company’s competitive advantage and substantial investments in data collection and training.
@@
1) To the best of our knowledge, we are the first to propose an algorithm that integrates foundation models into FL without distributing the foundation model to clients.
@@
When client data distributions are biased or non-IID [25], e.g., different ratios or a lack of certain labels, clients optimize corresponding local objective functions and may send conflicting updates to the server that skew the global model [26]. In FedBaF, the foundation model continuously serves as a form of regularization and stabilizes the global model by reducing the influence of these conflicting updates during the aggregation phase [3, 27, 28, 7].
Furthermore, FedBaF uses a fixed foundation model as an anchor and continuously incorporates it throughout the FL training process. This adds a layer of protection from adversarial attacks beyond those introduced by disclosing foundation model weights – such as misclassification or backdoor attacks, where compromised clients feed malicious updates to the server [11, 22].
3) We conduct extensive empirical evaluation and show that FedBaF matches or exceeds the training performance of traditional weight initialization methods – with better test performance in 10 out of 14 cases. Our experiments use Pre-ResNet and more complex architectures like Vision Transformer and Transformer-based language models frequently used as foundation models [29, 30]. Compared to standard FedAvg [10] and FedProx [31], FedBaF achieves accuracy improvements of up to 10.8% in IID and up to 37.5% globally and 5.9% locally in non-IID settings. Simultaneously, FedBaF safeguards the foundation model. Similarly, applying FedBaF to a Transformer-based language model significantly reduced perplexity by up to 76.0%.
Under adversarial backdoor misclassification attacks, FedBaF demonstrates increased robustness by improving FedAvg and FedProx test performance by up to 19.4% in IID environments, up to 64.7% globally, and 7.2% locally in non-IID environments. Additionally, in 8 out of 12 cases, FedBaF was more robust than traditional weight initialization methods.
@@ 2 Related Work
Traditionally, pre-trained models are used in FL to initialize the weights of the global FL model. The server distributes this model to local clients, and the clients update it with their local data. Such fine-tuning can significantly improve the FL model by integrating data from new clients, as supported by studies like FPS [3] and FedPCL [28].
Several works explore the use of foundation models in FL, particularly in scenarios with limited pre-training data [3]. This approach often involves using synthetic data [32, 3] for pre-training.
Federated Prototype-wise Contrastive Learning (FedPCL) is a significant development that aims to improve communication efficiency in FL by using class prototypes [28]. FedPCL enhances personalized learning by having clients share class-specific information more effectively. It fuses representations from multiple pre-trained models (backbones) at each client, with prototypes serving as the primary medium of information exchange. The strategy reduces communication overhead and enables customizing local models without extensive weight synchronization.
@@
The related works discussed in this section achieve good performance, but they do not consider the significant security vulnerabilities that result from sharing a foundation model with local clients, which compromise data privacy and the integrity of the global model. Malicious clients with access to foundation models can exploit them through: Model Inversion Attacks, recovering original training data or sensitive attributes from the model’s outputs [19, 21, 20]; Membership Inference Attacks, analyzing model predictions to determine whether specific data records were used in training [17, 16, 18]. These attacks compromise the security of the FL system, necessitating the development of more secure methods for leveraging pre-trained foundation models in FL settings.
@@
FedBaF addresses these security challenges by not sharing the foundation model with clients during the weight initialization stage. Instead, it dynamically integrates the foundation model’s pre-trained weights during the aggregation phase of each training round.
3 FedBaF Methodology
@@
In this section, we introduce FedBaF, whose approach is illustrated in Figure 2. During the aggregation phases of the FL training process, our FedBaF method repeatedly leverages pre-trained foundation model weights. For example, the pre-trained weights corresponding to feature extraction layers provide valuable representation mappings that guide the new model’s feature extractor during training. This approach can also extend to other layers if the foundation and FL models share the same architecture. When the new FL model has different neural network layers from the foundation model, i.e, the classifier layers, these different layers’ parameters are randomly initialized and then trained using data from the client pool, as in standard FL.
@@
Algorithm 1 describes how FedBaF fits into the traditional FL framework by incorporating foundation model weights during aggregation, as illustrated in Lines 9-11 [10]. FedBaF is versatile and can be embedded into many existing FL algorithms, e.g., SCAFFOLD [33], FedProx [31], or other FL strategies, by modifying their aggregation methods (Lines 8-11) and using their existing ClientUpdate$(\mathbf{w},D)$ logic for clients’ local training in Line 13.
@@
Modifying FL aggregation. FedBaF’s aggregation process in each training round begins with the same step as FedAvg’s or many other FL algorithms’ aggregation, by computing a weighted sum of the updated model parameters from each client (Line 8). After this aggregation, in Line 11 FedBaF incorporates the pre-trained model weights ($\mathbf{w}_{pre}$) into the global FL model, controlled by the factor $\tau_{t}$ defined in Line 9. Here, $(\mathbf{w}_{pre}\setminus\mathbf{w}_{t})$ refers to the subset of layers from the foundation model ($\mathbf{w}_{pre}$) that have the same architecture as the corresponding layers in the global FL model ($\mathbf{w}_{t}$), ensuring that only compatible layers are used during aggregation.
@@
Designing $\tau_{t}$. Our careful design of $\tau_{t}$ uses the $L2$ norm of the difference between consecutive normalized weights of $\mathbf{w^{\prime}}_{t+1}$ and $\mathbf{w}_{t}$, divided by $\sqrt{t+1}$. This change in the model’s weights between rounds reflects how much the global model adapts to new client updates. The normalization prevents $\tau_{t}$ from becoming too large. The factor $\sqrt{t+1}$ ensures that, as training progresses, the influence of $\mathbf{w}_{pre}$ gradually diminishes, but not too quickly, and $\mathbf{w}_{t+1}$ approaches the improving averaged weights $\mathbf{w^{\prime}}_{t+1}$. This strategy is critical to keeping the global model flexible and effective, especially when client data differs from the data used to train the foundation model [33].
@@
Depending on the network architectures (e.g., the number of weights or scale of the initialized weights), the scale of $\tau_{t}$ can vary. In non-IID situations and during adversarial attacks, the factor $\tau_{t}$ becomes significant. In particular, a large $\tau$ can indicate the presence of non-IID data or an attack, as such scenarios often result in large differences in consecutive weight updates. To keep $\tau_{t}$ within a suitable range, we introduce the parameter $\alpha_{t}$ in Line 10, which depends on $\tau_{0}$ and the hyper-parameter $\psi$. We empirically find that setting $\alpha_{t}$ such that $\alpha_{t}\tau_{0}$ is less than 2 in the initial round ($t=0$) prevents excessively large values that could overly bias the global model towards the foundation model. A lower bound of 1 for $\alpha_{t}\tau_{0}$ also ensures that the minimum impact of the foundation model is significant for small $t$. We thus ensure the influence of the foundation model in the critical initial training stages, while still allowing the global model to adapt as training progresses.
@@
In the worst-case scenario, where all local clients are malicious and collaborating to extract the foundation model’s weights, they can access $\tau_{t}$, $\mathbf{w}_{t+1}$, and $\mathbf{w^{\prime}}_{t+1}$. However, because $\alpha_{t}$ is randomly chosen in each round $t$ and known only to the server, the foundation model’s weights cannot be extracted. If $\alpha$ were static, even if the server did not disclose it, malicious clients could determine this constant value by solving the residual equations from two successive rounds,
$\frac{(1+\alpha\tau_{t})\mathbf{w}_{t+1}-\mathbf{w^{\prime}}_{t+1}}{\alpha\tau_{t}}=\frac{(1+\alpha\tau_{t+1})\mathbf{w}_{t+2}-\mathbf{w^{\prime}}_{t+2}}{\alpha\tau_{t+1}}.$
@@ 4 Theoretical Analysis of FedBaF
where $\mathbf{w}$ are model (usually, deep neural network) weights and the $f_{\mathcal{D}_{k}}$ are $L$-smooth local objective functions. The convergence analysis presented in this section also makes the following standard assumptions made by [33] and detailed in Appendix E. Each client locally optimizes $\mathbf{w}$ using stochastic gradient descent, where the stochastic gradients are (i) unbiased and (ii) have bounded variance. We also assume (iii) bounded gradient dissimilarity: the norm of the difference between the gradient of the global objective and the gradients computed using different local objective functions is bounded.

### v2

@@ 1 Introduction
Developing foundation models (Zhuang et al.,, 2023) has become a major focus for leading technology companies like OpenAI, Microsoft, and Amazon AWS. These deep learning models are often trained with vast amounts of high-quality data (Bommasani et al.,, 2021) and their ability to generalize across different tasks and domains has made them essential assets for industry, government, and academia. Foundation models have been applied to natural language processing (e.g., text generation, translation, summarization), image generation and recognition, healthcare diagnostics, finance predictive analytics, and customer service and virtual assistant tasks (Chen et al.,, 2023; Han et al.,, 2021; Duan et al.,, 2021; Joshi et al.,, 2022; Yosinski et al.,, 2014). When a foundation model’s training data distribution overlaps with a new application, it provides a robust starting point for fine-tuning and customization. Instead of training a model from scratch, with limited data and computes, we can leverage pre-trained foundation models to enable faster training.
For many applications, data that could be used to customize or fine-tune foundation models is often distributed across multiple clients, such as a network of clinics or small companies spread across different jurisdictions. For example, fine-tuning a recommendation model to fit a small company’s product offering may require data from clients in various regions; and adapting a healthcare model for a network of clinics would involve confidential, distributed data sources. Therefore, effective generalization requires access to diverse data from multiple clients (Pranav and Moura,, 2024).
Federated Learning (FL) is a promising solution for fine-tuning these models without sharing client data: FL clients train models on diverse local data, and a central FL server aggregates the client updates to build and refine a global model (Li et al.,, 2021; McMahan et al.,, 2017; Singh et al.,, 2019; Lyu et al.,, 2020; Nguyen et al.,, 2021; Wang et al.,, 2020; Lee et al.,, 2023; Siew et al.,, 2024). Using a foundation model to initialize the global FL model leads to effective customization that leverages diverse, distributed data without directly accessing the client data (Nguyen et al.,, 2022; Chen et al.,, 2023). However, there are significant risks associated with sending a foundation model to clients, as required in traditional FL fine-tuning methods.
@@
First, disclosing a foundation model’s weights to FL clients poses a significant security risk. For example, malicious actors could carry out membership inference attacks, identifying whether specific data was part of the foundation model’s training dataset. Then, an attacker could disrupt the global model’s training by introducing updates that degrade performance on identified data through backdoor attacks, deliberately leading to targeted misclassification (Dayal et al.,, 2023; Hu et al.,, 2022; Wang et al., 2023b, ). Similarly, for model inversion attacks, attackers use known model weights to reverse-engineer sensitive training data (see Figure 1(Fredrikson et al.,, 2015; Li et al., 2022a, ; Zhang et al.,, 2020). Protecting foundation models, often trained on sensitive, proprietary data, is critical for safeguarding the training data and maintaining model integrity (Bagdasaryan et al.,, 2020; Kim et al.,, 2023).
Second, in competitive business contexts, disclosing foundation model weights to clients risks leaking strategic insights and proprietary information to adversaries (Han et al.,, 2021; Yu et al.,, 2023). This undermines a company’s competitive advantage and substantial investments in data collection and training.
@@
When client data distributions are biased or non-IID, e.g., different ratios or a lack of certain labels, clients optimize correspondingly diverse local objective functions (Pranav and Moura,, 2023) and may send conflicting updates to the server that skew the global model (Zhao et al.,, 2018). In FedBaF, the foundation model continuously serves as a form of regularization and stabilizes the global model by reducing the influence of these conflicting updates during the aggregation phase (Chen et al.,, 2023; Li et al., 2022b, ; Tan et al.,, 2022; Yosinski et al.,, 2014).
Furthermore, FedBaF uses a fixed foundation model as an anchor and continuously incorporates it throughout the FL training process. This adds a layer of protection from adversarial attacks beyond those introduced by disclosing foundation model weights – such as misclassification attacks or backdoor attacks, where compromised clients feed malicious updates to the server (Lyu et al.,, 2020; Bagdasaryan et al.,, 2020).
@@
1)
To the best of our knowledge, we are the first to propose an algorithm that integrates foundation models into FL without distributing the foundation model to clients.
@@
3) We conduct extensive empirical evaluation and show that FedBaF matches or exceeds the training performance of traditional weight initialization methods – with better test performance in 10 out of 14 cases. Our experiments use Pre-ResNet and more complex architectures like Vision Transformer and Transformer-based language models frequently used as foundation models (Xu et al.,, 2023; Kenneweg et al.,, 2024). Compared to standard FedAvg (McMahan et al.,, 2017) and FedProx (Li et al.,, 2020) with foundation models used for weight initialization, FedBaF achieves accuracy improvements of up to 10.8% in IID and up to 37.5% globally and 5.9% locally in non-IID settings. Simultaneously, FedBaF safeguards the foundation model. Similarly, applying FedBaF to a Transformer-based language model significantly reduced perplexity by up to 76.0%.
Under adversarial misclassification attacks, FedBaF demonstrates increased robustness by improving FedAvg and FedProx test performance by up to 19.4% in IID environments, up to 64.7% globally, and 7.2% locally in non-IID environments. Additionally, in 8 out of 12 cases, FedBaF was more robust than traditional weight initialization methods.
@@ 2 Related Work
Traditionally, pre-trained models are used in FL to initialize the weights of the global FL model. The server distributes this model to local clients, and the clients update it by using their local data. We refer to this approach as “weight initialization” throughout this paper. Such fine-tuning of a pre-trained model can significantly improve performance of the learned FL global model by integrating data from new clients (Nguyen et al.,, 2022). Several recent studies devised methods that leverage weight initialization to further improve performance: Federated Nearest Class Means (FedNCM) (Legate et al.,, 2023) for last-layer guidance, Federated Recursive Ridge Regression (Fed3R) (Fanì et al.,, 2024), Fractal Pair Similarity (FPS) (Chen et al.,, 2023), and FedPCL (Tan et al.,, 2022).
Several works also explore the use of foundation models in FL. These include cases where a subset of the weights of a large foundation model are chosen to initialize and fine-tune a smaller model (Xu et al.,, 2024) and when clients have diverse model architectures (Wang et al., 2023a, ; Park and Joe-Wong,, 2024). Particularly in scenarios with limited pre-training data (Chen et al.,, 2023), approaches often rely on synthetic data (Nikolenko,, 2021; Chen et al.,, 2023) for pre-training. Federated Prototype-wise Contrastive Learning (FedPCL) is a significant development that improves communication efficiency in FL by using class prototypes (Tan et al.,, 2022) and enhances personalized learning by having clients share class-specific information more effectively.
@@
The related works discussed in this section so far achieve good performance, but they do not consider the significant security vulnerabilities that result from sharing a foundation model with local clients, which compromise data privacy and the integrity of the global model. Malicious clients with access to foundation models can exploit them through: Model Inversion Attacks, recovering original training data or sensitive attributes from the model’s outputs (Fredrikson et al.,, 2015; Zhang et al.,, 2020; Li et al., 2022a, ); Membership Inference Attacks, analyzing model predictions to determine whether specific data records were used in training (Hu et al.,, 2022; Dayal et al.,, 2023; Wang et al., 2023b, ). These attacks compromise the security of the FL system, necessitating the development of more secure methods for leveraging pre-trained foundation models in FL settings.
@@
FedBaF addresses these security challenges by not sharing the foundation model with clients during the weight initialization stage. Instead, it dynamically integrates the foundation model’s pre-trained weights during the aggregation phase of each training round. We show that FedBaF improves privacy while matching or exceeding the performance achieved by weight initialization.
3 Methodology
@@
In this section, we introduce FedBaF, whose approach is illustrated in Figure 2. FedBaF involves the server repeatedly leveraging pre-trained foundation model weights throughout the aggregation phases of the FL training process. For example, the pre-trained weights corresponding to feature extraction layers provide valuable representation mappings that guide the new model’s feature extractor during training.
To mimic the performance gains of weight initialization, the server uses the foundation model as a strong anchor in the earlier FL rounds. To enable the FL global model to evolve and fit the clients’ data as FL training continues, the server assigns rapidly decaying importance to the foundation model that is on the order of $1/\sqrt{t}$ and proportional to the change in model parameters caused by client updates. To further maintain foundation model confidentiality, the server randomly samples the aggregation weights (or importance) of the foundation model from a uniform distribution in each round. These aspects of FedBaF maintain foundation model privacy while enabling further application-specific tuning to achieve performance on par or exceeding that of weight initialization methods.
@@
Algorithm 1 describes how FedBaF fits into the traditional FL framework by incorporating foundation model weights during aggregation, as illustrated in Lines 9-11 (McMahan et al.,, 2017). FedBaF is versatile and can be embedded into many existing FL algorithms, e.g., SCAFFOLD (Karimireddy et al.,, 2020), FedProx (Li et al.,, 2020), FedAdam (Reddi et al.,, 2021), or other FL strategies, by modifying their aggregation methods (Lines 8-11) and using their existing ClientUpdate$(\mathbf{w},D)$ logic for clients’ local training in Line 13.
@@
Modifying FL aggregation. FedBaF’s aggregation process in each training round begins with the aggregation step of an existing FL algorithm, which, as mentioned above, includes FedAvg, SCAFFOLD, FedProx, and FedAdam. To illustrate an example with FedAvg, Line 8 of Alg. 1 uses FedAvg’s aggregation step and computes a weighted sum of the updated model parameters from each client. After this aggregation, Line 11 incorporates the pre-trained model weights ($\mathbf{w}_{pre}$) into the global FL model, controlled by the factor $\tau_{t}$ defined in Line 9.
Here, $(\mathbf{w}_{pre}\setminus\mathbf{w}_{t})$ refers to the subset of layers from the foundation model ($\mathbf{w}_{pre}$) that have the same architecture as the corresponding layers in the global FL model ($\mathbf{w}_{t}$), ensuring that only compatible layers are used during aggregation. When the foundation and FL models have the same architecture except that the input and output layers (i.e., first and last layers) differ due to variations in input features or the number of classes, FedBaF excludes these input and output layers from the aggregation and integrates only the shared intermediate (hidden) layers. The differing layers are randomly initialized and then trained using data from the client pool, as in standard FL.
Foundation and Global Model Architecture Mismatch. In practice, the foundation model and the global model architectures may differ beyond the input and output layers. These differences can arise in terms of the number of layers or the number of parameters per layer. Since foundation models are typically larger than global FL models – consistent with their role as highly expressive networks pre-trained on extensive data (Meng et al.,, 2023; Awais et al.,, 2025) – we consider the following two cases:
$\bullet$ Foundation model with more layers: When the foundation model is deeper than the global FL model, only a subset of its layers is used during aggregation. We take advantage of the fact that most large models have layers grouped into sections. Here, a section is a contiguous subset of layers within a model that shares similar structural properties, such as the number of parameters, functional roles, or connectivity patterns. FedBaF selects and matches sections between the foundation and global models, prioritizing feature extraction sections near the input layer. Within each matched section, only layers that align with the global model’s architecture are integrated into FL training. Methods for selecting compatible layers during aggregation are explored in Park and Joe-Wong, (2024); Xu et al., (2024).
$\bullet$ Foundation model with more parameters per layer: If the foundation model has layers with more parameters than the global model, only a subset of parameters within each section and layer is selected to match the global FL model. Since different sections may have varying parameter distributions, the aggregation is performed iteratively per section and per layer. To consider these mismatches, Line 11 of Alg. 1 is expanded:
1:  for each section $s$ in shared sections between $\mathbf{w}_{pre}$ and $\mathbf{w^{\prime}}_{t}$ do
2:     for each layer $l$ in section $s$ do
3:        for each parameter subset $p$ in layer $l$ up to $P_{t}^{(s,l)}$ parameters do
4:           $\mathbf{w}_{t+1}^{(s,l,p)}\leftarrow\frac{1}{1+\alpha_{t}\tau_{t}}\left(\mathbf{w^{\prime}}_{t+1}^{(s,l,p)}+\alpha_{t}\tau_{t}\mathbf{w}_{pre}^{(s,l,p)}\right)$
5:        end for
6:     end for
7:  end for
Here, $P_{t}^{(s,l)}$ denotes the number of parameters selected per layer $l$ within section $s$. Only the first $P_{t}^{(s,l)}$ parameters in each layer are incorporated into the global model.
FedBaF can be analogously modified to handle the case where the foundation model is smaller than the global FL model. By applying these modifications, FedBaF can seamlessly adapt to different network architectures, ensuring that foundation model knowledge is effectively transferred while maintaining structural compatibility with the FL model.
@@
Designing $\tau_{t}$. Our careful design of $\tau_{t}$ uses the $L2$ norm of the difference between consecutive normalized weights of $\mathbf{w^{\prime}}_{t+1}$ and $\mathbf{w}_{t}$, divided by $\sqrt{t+1}$. This change in the model’s weights between rounds reflects how much the global model adapts to new client updates. The normalization prevents $\tau_{t}$ from becoming too large. The factor $\sqrt{t+1}$ ensures that, as training progresses, the influence of $\mathbf{w}_{pre}$ gradually diminishes, but not too quickly, and $\mathbf{w}_{t+1}$ approaches the improving averaged weights $\mathbf{w^{\prime}}_{t+1}$. This strategy is critical to keeping the global model flexible and effective, especially when client data differs from the data used to train the foundation model (Karimireddy et al.,, 2020).
@@
Depending on the network architectures (e.g., the number of weights or scale of the initialized weights), the scale of $\tau_{t}$ can vary. In non-IID situations and during adversarial attacks, the factor $\tau_{t}$ becomes significant. In particular, a large $\tau_{t}$ can indicate the presence of non-IID data or an attack, as such scenarios often result in large differences in consecutive weight updates. To keep $\tau_{t}$ within a suitable range, we introduce the parameter $\alpha_{t}$ in Line 10, which depends on $\tau_{0}$ and the hyper-parameter $\psi$. We empirically find that setting $\alpha_{t}$ such that $\alpha_{t}\tau_{0}$ is less than 2 in the initial round ($t=0$) prevents excessively large values that could overly bias the global model towards the foundation model. A lower bound of 1 for $\alpha_{t}\tau_{0}$ also ensures that the minimum impact of the foundation model is significant for small $t$. We thus ensure the influence of the foundation model in the critical initial training stages, while still allowing the global model to adapt as training progresses.
@@
In the worst-case scenario, where all local clients are malicious and collaborating to extract the foundation model’s weights, they can access $\tau_{t}$, $\mathbf{w}_{t+1}$, and $\mathbf{w^{\prime}}_{t+1}$. However, because $\alpha_{t}$ is randomly chosen in each round $t$ and known only to the server, the foundation model’s weights cannot be extracted. If $\alpha_{t}$ were static, even if the server did not disclose it, malicious clients could determine this constant value by solving the residual equations from two successive rounds,
$\frac{(1+\alpha_{t}\tau_{t})\mathbf{w}_{t+1}-\mathbf{w^{\prime}}_{t+1}}{\alpha_{t}\tau_{t}}=\frac{(1+\alpha_{t+1}\tau_{t+1})\mathbf{w}_{t+2}-\mathbf{w^{\prime}}_{t+2}}{\alpha_{t+1}\tau_{t+1}}.$
@@ 4 THEORETICAL ANALYSIS
where $\mathbf{w}$ are model (usually, deep neural network) weights and the $f_{\mathcal{D}_{k}}$ are $L$-smooth local objective functions. The convergence analysis presented in this section also makes the following standard assumptions made by (Karimireddy et al.,, 2020) and detailed in Appendix E. Each client locally optimizes $\mathbf{w}$ using stochastic gradient descent, where the stochastic gradients are (i) unbiased and (ii) have bounded variance. We also assume (iii) bounded gradient dissimilarity: the norm of the difference between the gradient of the global objective and the gradients computed using different local objective functions is bounded. Lastly, we assume that (iv) the foundation model has the same architecture as the global model, ensuring compatibility during aggregation. See Appendix E.1.1 for mathematical details regarding the assumptions.